This year, the Tastee Partnership reported income before guaranteed payments of $196,500. Stella owns a 50% profits interest and works 1,730 hours per year in the business. Euclid owns a 50% profits interest (with a basis of $30,000 at the beginning of the tax year) and performs no services for the partnership during the year. For services performed during the year, Stella receives a "salary" of $9,825 per month. Euclid withdrew $19,650 from the partnership during the year.

a. What is the amount of guaranteed payments made by the partnership this year? $

b. How much is the partnership's ordinary income after any permitted deduction for guaranteed payments? $

c. How much income will Stella and Euclid report? Stella: $ Euclid: $
Answer #1
Guaranteed Payment= "Guaranteed Payments are those payments made by partnership to a partner that is determined without regard to partnership income."
a) Guaranteed Payment made by partnership this year=Stella's Salary=($9825*12)= $                                  1,17,900.00
b) Partnership's ordinary income after any permitted deduction for guaranteed payment=Partnership reported income-Stella's Salary
Partnership reported income=(A) $                                  1,96,500.00
Stella's Salary=(B) $                                  1,17,900.00
Partnership's ordinary income after any permitted deduction for guaranteed payment=(A)-(B) $                                      78,600.00
c ) Stella's Income=(50% of Ordinary Income after guaranteed payment+Guaranteed Payment)
90% of Ordinary Income=($78600*50%)=(A) $                                      39,300.00
Guaranteed Payment=(B) $                                  1,17,900.00
Stella's Income=(A)+(B) $                                  1,57,200.00
Euclid's Income=(50% of ordinary income after guaranteed payment)=($78600*50%) $                                      39,300.00